the length of a rectangle is 5 more than twice it's width. its perimeter is 88 feet. find it's dimensions
Answer by Alan3354(69443)

2L + 2W = 88
L + W = 44
L = 2W + 5
----------
Sub for L into line 2
(2W+5) + W = 44
3W = 39
W = 13
L = 31
------
PS it's = it is.
the length of a rectangle is 5 more than twice its width. its perimeter is 88 feet. find its dimensions